1. What is the representation of -9 in ten’s complement using two decimal digits?


2. What is the representation of -9 in ten’s complement using three decimal digits?


3. Which of the following is represented as 11111100 in two’s complement ?

A. -2

B. -4

C. -6

D. 252

E. -252
Answer #1

1) 10s complement means 9s complement of a decimal no. + adding 1 to least significant bit

so for -9 ............. the 10s complement up to 2 decimals ...........9.99-(-9)=18.99 is 9s complement .

now add 0.01 to obtained answer 18.99+0.01=19.00

2) 10s   complement up to 3 decimals ...........for -9 is 9.999-(-9)=18.999 is 9s complement.

now add 0.001 to obtained answer 18.999 +0.001 =19.000

3) 11111100 is in 2s complement

so first do 1s complement to this(11111100).........is 00000011 and now add 1 to gets its original number ..

0 0 0 0 0 0 1 1

+ 1

0 0 0 0 0 1 0 0

so we got 00000100 as the answer so converting it to decimal as 0*20+0*21+1*22+0+0+0+0+0=4 as its a negative number we indicate it with negative sign

so answer is -4